# Universal Post-Training Reverse-Engineering Defense Against Backdoors in Deep Neural Networks

### v1

@@ 3.1 Discussion: Hyperparameters
One might consider the choice of the internal layer $f(\cdot)$ and the Lagrange multiplier $\lambda$ to be hyperparameters of our detector. However, our detection approach can be applied to various layers of the DNN, with detection inference then based only on the layer with the most extreme detection statistic (smallest p-value). Thus, the internal layer can be chosen in an automated fashion (and is not really a hyperparameter of the algorithm). Similarly, as discussed next in the experiments section, $\lambda$ can also be automatically adjusted, and is thus also not truly a hyperparameter of the algorithm. The threshold on the misclassification rate (used to terminate the CEPA optimization) is a hyperparameter of the algorithm. However, we have found experimentally that CEPA’s performance is largely insensitive to the choice of this threshold.
@@ 4.1 Experiment Setup
Datasets: Our experiments are conducted on the benchmark CIFAR-10 dataset, which contains 60,000 $32\times 32$ color images from 10 classes, with 5,000 images per class for training and 1,000 images per class for testing [10].
Attack Settings: On CIFAR-10, we consider the following triggers: 1) a $3\times 3$ random patch (BadNet) with a randomly selected location (fixed for all triggered images for each attack) used in [6]; 2) an additive perturbation (with size 2/255) resembling a chessboard used in [22]; 3) a warping-based trigger (WaNet) proposed by [14]; 4) a Hello Kitty trigger with a blend ratio of $\alpha=0.15$ used by [4]. We mainly considered all-to-one attacks, and arbitrarily chose class 9 as the target class. We set the poisoning rate at 1.8%. That is, we embedded the backdoor triggers in 100 randomly chosen training samples per class (excluding the target class). Moreover, to achieve similarly effective attacks as for other triggers, we poisoned 900 images per source class in the all-to-one attack for WaNet.
Training Settings: We trained a ResNet-18 ([7]) on CIFAR-10 for 30 epochs with a batch size of 32. For all attacks, we used the Adam optimizer ([9]) for parameter learning and a scheduler to decay the learning rate of each parameter group by 0.1 every 10 epochs. The initial learning rate was set to 0.01.
Defense Settings: For trigger estimation, we randomly selected 10 images per class from the test set to form the small clean dataset possessed by the defender. The remaining test instances were reserved for performance evaluation. For each putative target class, the parameter $\lambda$ of (3) was initially set to $0.01$ and dynamically adjusted based on the misclassification rate: If it is higher than 90% for 5 consecutive iterations, $\lambda$ is increased by a factor of 1.5 (tightening the embedded consensus around the current $\mu$). If it is lower than 90% for 5 consecutive iterations, $\lambda$ is decreased by a factor of 1.5. The optimization of (3) was terminated if the mean norm of the $\delta$s did not decrease over 50 consecutive iterations while the misclassification rate is maintained over 90%.
@@ 4.2 Backdoor Detection Performance
Acting on layer 5 of the ResNet18 model, Figures 1 and 2 respectively show the performance of our defense in terms of the consensus measure consisting of the estimated standard deviation divided by the L2 norm of $\mu$, $\sigma/\|\mu\|$, and in terms of $\|\mu\|$, on the four poisoned models and on an unattacked (clean) model. Figures 3 and 4 show the same measures for layer 13. From the bar charts (one bar for each putative target class), one can see that, for both layers 5 and 13, our method strongly indicates the true target class (9) on poisoned models, i.e., the quantities for the true target class are clear outliers with respect to the non-target classes (0-8) [^2] Such outlier detection is directly accomplished by applying, e.g., Median Absolute Deviation (MAD), on the null formed by the bar chart, as used in Neural Cleanse [18]., but does not indicate any significant outliers in the case of a clean model (i.e., there are no false positive class detections). The exception is the global chessboard backdoor, which is only clearly detected using layer 5. We again note that our method can be applied to a plurality of layers (embedded feature representations) of the model to detect backdoor poisoning.
@@
Fig. 1: Layer-5 consensus, $\sigma/\|\mu\|$
@@
Fig. 3: Layer-13 consensus, $\sigma/\|\mu\|$
@@
Fig. 4: Layer-13 L2 norm of $\mu$, $\|\mu\|$
Fig. 5: Visualization of the BadNet attack: (a) image with the ground truth trigger, (b) image with the estimated trigger.
(a) chessboard
(b) BadNet
Fig. 6: Mean pairwise cosine similarity of the optimizing $\delta$s.
Some prior work on backdoor detection has suggested the use of cosine similarities of sample-specific perturbations $\delta$ as a detection statistic. In Figure 6, we plot the mean (pairwise) cosine similarity of the optimizing $\delta$s obtained from our method for each putative target class. Though the mean cosine similarity is relatively larger for the backdoor target class for chessboard than for BadNet (chessboard is an additive perturbation (which is not sample-specific), while BadNet’s patch is sample-specific because it involves replacing the portion of the image in the backdoor mask), the mean cosine similarity is not effective as a detection statistic, for either of the two attacks – this shows that $\|\mu\|$ and the consensus measure $\sigma/\|\mu\|$ (inspired by the backdoor overfitting phenomenon recognized in [19]) are quite distinctive as (sound) backdoor detection statistics.
ASR patch blend WaNet chessboard
@@
Table 1: Attack success rates (for a 1.8% poisoning fraction of the training set) when: i) into the test samples, the ground-truth (GT) backdoor pattern is inserted; ii) perturbations $\delta_{x}$ obtained from (3) are inserted into the small clean dataset used to find $\mu$; and iii) perturbations obtained via (3) (given $\mu$) are inserted into the test samples.
In Table 1, attack success rates (ASRs) are evaluated on the test set when the ground-truth backdoor pattern is inserted into the test samples and when, instead, the perturbations obtained via (3) are inserted into test samples. (We also show the ASR on the clean set when the estimated $\delta$s are inserted into the clean set samples.) Specifically, we
  • randomly chose 100 correctly classified samples per (detected) source class from the CIFAR-10 test dataset,
@@
    $\delta_{x}=\operatorname*{arg\,min}_{\delta}\|f(x+\delta)-(f(x)+\mu)\|^{2}+\upsilon\|\delta\|^{2},$ (3)
@@
While UNICORN [21] reports higher ASRs (see their Table 1), their poisoning fraction is $2.8$ times higher (5%, see their Section A.5, compared to our 1.8%), which may make the backdoor pattern more conspicuous and more easily detected. Note, though, that a poisoning rate of just 1.8% is sufficient to achieve a very high attack success rate for the adversary (Table 1). As mentioned above, if UNICORN’s feature maps are trained using the (possibly poisoned) training dataset of the defended model, then it is not truly a post-training defense. Again, the post-training backdoor detection scenario is important considering that, e.g., AIs are sometimes built from foundation models acquired from other parties or are produced by other parties and embedded in larger systems which are then acquired (with potentially no “data rights” access to the training set, for the party who purchases/acquires the model). Additionally, CEPA
@@
  • does not, in particular, involve a bound on the “size” of the backdoor pattern or its mask (unlike UNICORN’s “mask size constraint”) – indeed, it detects the chessboard pattern, which has global pixel support;
@@ 4.3 Estimated Pattern Visualization and Analysis
Visualizing Estimated Triggers. In Figure 5, we visualize compare the estimated backdoor pattern/trigger for the BadNet attack using CEPA with the ground-truth trigger. Specifically, Figure 5 (a) displays a clean image from class 0 (airplane) embedded with the ground truth trigger (lower right of the image). Figure 5 (b) is the same clean image embedded with the reverse-engineered ($\delta_{x}$) trigger.
Note that, in CEPA’s loss function, triggers are (sample-wise) additively applied to images. The resulting, estimated sample-wise $\delta_{x}$’s may in fact have non-zero values at all pixel locations (albeit small values at pixel locations that are not part of the ground-truth backdoor pattern’s support, for localized patch replacement attacks). A notable observation from these visualizations is the discrepancy in appearance between the ground truth and the estimated trigger, despite their shared location in the image’s right lower corner. The similarity in positioning highlights the precision of our estimation process in identifying the portion of the image that has been adversarially manipulated. However, the visual differences between the two triggers underscores that is challenging to perfectly replicate adversarial patterns (reflecting the complexity of such cyber-threats and the need for nuanced approaches for analyzing and mitigating them).
@@ 5 Summary and Discussion
This paper concerns post-training defense against backdoor poisoning in deep neural networks. The proposed method CEPA detects whether the model is backdoor poisoned and, if so, identifies the target class of the backdoor and also estimates the backdoor trigger pattern for a given sample. CEPA is agnostic to the manner of incorporation of the backdoor pattern (as well as agnostic to the backdoor pattern itself, of course). CEPA has low complexity and so is scalable (even when based on a plurality of embedded feature representations); requires few clean samples; requires very few hyperparameters (effectively, only a single one); relies only on embedded feature activations (not on auxiliary feature maps); and performs well, with our results easily reproduced by other parties. As noted earlier, $\lambda$ can be dynamically adjusted during the optimization process, while different choices for the layer index of the embedded feature representation can be simultaneously assessed without large computational overhead. CEPA detection is also designed to work when there are multiple backdoor attacks present with different target classes, so long as their number is significantly smaller than the total number of classes.

### v2

@@ 3.1 Discussion: Hyperparameters
One might consider the choice of the internal layer $f(\cdot)$ and the Lagrange multiplier $\lambda$ to be hyperparameters of our detector. However, our detection approach can be simultaneously applied to various layers of the DNN, with detection inference then based only on the layer with the most extreme detection statistic (smallest p-value). Thus, the internal layer can be chosen in an automated fashion (and is not really a hyperparameter of the algorithm). Similarly, as discussed next in the experiments section, $\lambda$ can also be automatically adjusted, and is thus also not truly a hyperparameter of the algorithm. The threshold on the misclassification rate (used to terminate the CEPA optimization) is a hyperparameter of the algorithm. However, we have found experimentally that CEPA’s performance is largely insensitive to the choice of this threshold.
@@ 4.1 Experiment Setup
Datasets: Our experiments are conducted on the benchmark datasets CIFAR-10 and CIFAR-100 [10]. CIFAR-10 contains 60,000 $32\times 32$ color images from 10 classes, with 5,000 images per class for training and 1,000 images per class for testing . CIFAR-100 contains 60,000 $32\times 32$ color images from 100 classes, with 500 images per class for training and 100 images per class for testing.
Attack Settings: On CIFAR-10, we consider the following triggers: 1) a $3\times 3$ random patch (BadNet) with a randomly selected location (fixed for all triggered images for each attack), used in [6]; 2) an additive perturbation (with size 2/255) resembling a chessboard [24]; 3) a warping-based trigger (WaNet) proposed by [14]; 4) a Hello Kitty trigger with a blend ratio of $\alpha=0.15$ used by [4]. We mainly considered all-to-one attacks, and arbitrarily chose class 9 as the target class. We set the poisoning rate at 1.8%. That is, we embedded the backdoor triggers in 100 randomly chosen training samples per class (excluding the target class). Moreover, to achieve similarly effective attacks as for other triggers, we poisoned 900 images per source class in the all-to-one attack for WaNet. Our CIFAR-100 experiment illustrates the scalability of CEPA. On CIFAR-100 we only consider the BadNet attack.
Training Settings: We trained a ResNet-18 ([7]) on CIFAR-10 and a VGG-16 ([16] on CIFAR-100 for 30 epochs with a batch size of 32. For all attacks, we used the Adam optimizer ([9]) for parameter learning and a scheduler to decay the learning rate of each parameter group by 0.1 every 10 epochs. The initial learning rate was set to 0.01.
Defense Settings: For trigger estimation, we randomly selected 10 images per class from the test set to form the small clean dataset possessed by the defender. The remaining test instances were reserved for performance evaluation. For each putative target class, the parameter $\lambda$ of (3) was initially set to $0.01$ and dynamically adjusted based on the misclassification rate: If it is higher than 90% for 5 consecutive iterations, $\lambda$ is increased by a factor of 1.5 (tightening the embedded consensus around the current $\mu$). If it is lower than 90% for 5 consecutive iterations, $\lambda$ is decreased by a factor of 1.5. The minimization of (3) is terminated when the mean norm of the $\delta$s ($|\mathcal{D}_{-t}|^{-1}\sum_{x\in\mathcal{D}_{-t}}\|\delta_{x}\|$ where $\mathcal{D}_{-t}=\cup_{s\not=t}\mathcal{D}_{s}$) did not decrease over 50 consecutive iterations while the misclassification rate is maintained over 90%.
@@ 4.2 Backdoor Detection Performance
Acting on layer 5 of the ResNet18 model, Figures 1 and 2 respectively show the performance of our defense in terms of the consensus measure consisting of the estimated standard deviation divided by the L2 norm of $\mu$, $\sigma/\|\mu\|$, and in terms of $\|\mu\|$, on the four poisoned models and on an unattacked (clean) model. Figures 3 and 4 show the same measures for layer 13. From the bar charts (one bar for each putative target class), one can see that, for both layers 5 and 13, our method strongly indicates the true target class (9) on poisoned models, i.e., the quantities for the true target class are clear outliers with respect to those for the non-target classes (0-8) [^4] Such outlier detection is directly accomplished by applying, e.g., Median Absolute Deviation (MAD), on the null formed by the bar chart, as used in Neural Cleanse [19]., but does not indicate any significant outliers in the case of a clean model (i.e., there are no false positive class detections). The exception is the global chessboard backdoor, which is only clearly detected using layer 5. We again note that our method can be applied to a plurality of layers (embedded feature representations) of the model to detect backdoor poisoning.
Table 1 shows CEPA detection’s MAD statistics, e.g., [19] [^5] Order-statistic p-values could be used instead of MAD, e.g., [24].. Note that the chessboard pattern is not detected (i.e., MAD $<2$, where 2 is the traditional detection threshold corresponding to two standard deviations beyond the median) using layer 13 with both L2 norm of $\mu$ and consensus, but is detected at layer 5. Also, the clean model is “borderline" detected at layer 13 but not detected at layer 5. Again, CEPA can be run using a plurality of layers and both criteria, with a detection made, e.g., if any corresponding MAD in a range $\geq 2$. The MAD anomaly scores for all classes and all attacks are based on the data shown in Figures 1, 2, 3, and 4. Also from these bar charts (one bar for each putative target class), one can see that, for both layers 5 and 13, CEPA strongly indicates the true target class (9) on poisoned models, i.e., the quantities for the true target class are clear outliers with respect those for the non-target classes (0-8), but does not indicate any significant outliers in the case of a clean model (i.e., there are no false positive class detections). The exception is the global chessboard backdoor, which is only clearly detected using layer 5.
Table 1: For each DNN layer and detection criterion, CEPA MAD statistics for the attacks against ResNet-18/CIFAR-10, and also for the Clean (unpoisoned) model.
Layer and criterion BadNet WaNet Chessboard Blend Clean
5, L2 norm of $\|\mu\|$ 27.47 7.11 7.48 7.25 0.50
5, consensus ($\sigma/\|\mu\|$) 2.54 4.12 5.72 2.25 0.23
13, L2 norm of $\|\mu\|$ 9.13 4.10 0.23 3.91 0.75
13, consensus ($\sigma/\|\mu\|$) 3.58 2.72 0.34 1.71 2.02
@@
Figure 1: Layer-5 consensus, $\sigma/\|\mu\|$
@@
Figure 3: Layer-13 consensus, $\sigma/\|\mu\|$
@@
Figure 4: Layer-13 L2 norm of $\mu$, $\|\mu\|$
Figure 5: For the BadNet attack: (a) an image $x$ with the ground truth trigger, (b) an image $x$ with its estimated trigger $\hat{\delta}_{x}$, (c) $|\mathcal{D}|^{-1}\sum_{x\in\mathcal{D}}\hat{\delta}_{x}$ with pixel intensities amplified by 5 for visualization.
(d) chessboard
(e) BadNet
Figure 6: Mean pairwise cosine similarity of the optimizing $\delta$s.
Figure 7: Histogram of MAD anomaly scores calculated on (a) L2 norm of $\mu$ and (b) consensus in layer-5, (c) L2 norm of $\mu$ and (d) consensus in layer-10.
Some prior work on backdoor detection has suggested the use of cosine similarity as a detection statistic, e.g., [21]. In Figure 6, we plot the mean (pairwise) cosine similarity of the optimizing $\delta$s obtained from our method for each putative target class. Though the mean cosine similarity is relatively larger for the backdoor target class for chessboard than for BadNet (chessboard is an additive perturbation (which is not sample-specific), while BadNet’s patch is sample-specific because it involves replacing the portion of the image in the backdoor mask), the mean cosine similarity is not effective as a detection statistic, for either of the two attacks – this shows that $\|\mu\|$ and the consensus measure $\sigma/\|\mu\|$ (inspired by the backdoor overfitting phenomenon recognized in [20]) are quite distinctive as (sound) backdoor detection statistics.
ASR BadNet Blend WaNet chessboard
@@
Table 2: Attack success rates (for a 1.8% poisoning fraction of the training set) when: i) the ground-truth (GT) backdoor pattern is incorporated into the evaluation samples; ii) perturbations $\delta$ obtained from (3) are incorporated into the samples of the small clean dataset $\mathcal{D}$ use for defense; and iii) perturbations $\delta$ obtained via (3) (given $\mu$ from (3)) are incorporated into the evaluation set.
In Table 2, attack success rates (ASRs) are evaluated on the test set when the ground-truth backdoor pattern is added to the test samples and when, instead, the perturbations obtained via (3) are added to test samples. Specifically, for evaluation, we
  • randomly chose 100 correctly classified samples per source class from the CIFAR-10 test dataset,
@@
    $\hat{\delta}_{x}=\operatorname*{arg\,min}_{\delta}\|f(x+\delta)-(f(x)+\mu)\|^{2}+\upsilon\|\delta\|^{2},$ (3)
@@
From Table 2, we see that, with the exception of chessboard, using reverse-engineered perturbations based on $\mu$ (3) on evaluation examples results in high ASR but not as high as using either the ground-truth (GT) perturbations nor the reverse-engineered perturbations on the samples used by the defense to identify $\mu$ (3). It is interesting that in one case (blend), the GT ASR is lower than that of (3).
While UNICORN [23] reports higher ASRs (see their Table 1), their poisoning fraction is $2.8$ times higher (5%, see their Section A.5, compared to our 1.8%), which may make the backdoor pattern more conspicuous and more easily detected. Note, though, that a poisoning rate of just 1.8% is sufficient to achieve a very high attack success rate for the adversary (Table 2).
As mentioned above, if UNICORN’s feature maps are trained using the (possibly poisoned) training dataset of the defended model, then it is not truly a post-training defense. Again, the post-training backdoor detection scenario is important considering that, e.g., AIs are sometimes built from foundation models acquired from other parties or are produced by other parties and embedded in larger systems which are then acquired (with potentially no “data rights” access to the training set, for the party who purchases/acquires the model).
Additionally, CEPA
@@
  • does not involve a bound on the “size" of the backdoor pattern or its mask (unlike UNICORN’s “mask size constraint") – indeed, it detects the chessboard pattern, which has global pixel support;
@@
CEPA has similarly good detection results and scalability on the 5th layer of VGG-16 model trained on CIFAR-100, as shown in Figure 7 against BadNet. More details are in Appendix A.
Also, see Appendix B for results of an adaptive BadNet attack on a ResNet-18 model designed for CIFAR-10.
@@ 4.3 Estimated Pattern Visualization and Analysis
Visualizing Estimated Triggers. In Figure 6, we visually compare the estimated backdoor pattern/trigger for the BadNet attack using CEPA with the ground-truth trigger: (a) is a clean image from class 0 (airplane) embedded with the ground truth trigger (lower right of the image); (b) is the same clean image embedded with the reverse-engineered ($\delta_{x}$) trigger; and (c) is the average estimated $\hat{\delta}$ over all clean samples of the source class, $\mathcal{D}_{s}$.
Note that, in CEPA’s loss function, triggers are (sample-wise) additively applied to images. The resulting, estimated sample-wise $\delta_{x}$’s may in fact have non-zero values at all pixel locations (albeit small values at pixel locations that are not part of the ground-truth backdoor pattern’s support, for localized patch replacement attacks).
A notable observation from these visualizations is that, while there is discrepancy in appearance between the ground truth and the estimated learned triggers, they are in the same location.
4.4 Time Complexity
Our experiments utilize an NVIDIA RTX 3090 GPU with 24GB of memory. CEPA requires approximately 2 hours to perform detection on a Resnet-18 model trained on CIFAR-10, using 10 clean images per class (90 images total). Conversely, detection on a VGG-16 model trained on CIFAR-100 using 2 clean images per class (198 images total) takes approximately 40 hours. For comparison, training a Resnet-18 on CIFAR-10 for 30 epochs takes approximately 20 minutes.
@@ 5 Summary and Discussion
This paper concerns post-training defense against backdoor poisoning in deep neural networks. The proposed method CEPA detects whether the model is backdoor poisoned and, if so, identifies the target class of the backdoor and also estimates the backdoor trigger pattern for a given sample. CEPA is agnostic to the manner of incorporation of the backdoor pattern (as well as agnostic to the backdoor pattern itself, of course). CEPA has low complexity and so is scalable (even when based on a plurality of embedded feature representations); requires few clean samples; requires very few hyperparameters (effectively, only a single one); relies only on embedded feature activations (not on auxiliary feature maps); and performs well, with our results easily reproduced by other parties. As noted earlier, $\lambda$ can be dynamically adjusted during the optimization process, while different choices for the layer index of the embedded feature representation can be simultaneously assessed without large computational overhead.
Appendix A BadNet against VGG-16 for CIFAR-100
VGG has fewer (fatter) layers than the comparable ResNet model, the layers selected to apply CEPA are more sensitive hyperparameters. Note that if we choose layers too close to the input, we don’t get a suitable feature map for the embedded activation of the backdoor to manifest. Alternatively, if we choose a layer too close to the output, the target class "logit" signal dominates for the backdoor trigger.
Table 3: CEPA MAD statistics against BadNet on VGG-16/CIFAR-100
[…]
5 28.44 14.03
10 0.52 0.51
In Table 3, we show CEPA detection’s MAD statistics for the poisoned model. Both are $>2$, i.e., the backdoor is successfully detected, at layer 5. Also, just as for the CIFAR-10 results in Table 1, there were no false positives for the clean/unpoisoned CIFAR-100 model.
Appendix B Adaptive BadNet attack on ResNet-18 for CIFAR-10
Assume the attacker is the training authority and trains the model to minimize consensus (maximize variance of embedded perturbation corresponding to the backdoor). That is, for the poisoned training data subset $\mathcal{X}_{\rm B}$ of the training dataset $\mathcal{X}$, the training objective would be to minimize the cross-entropy loss over $\mathcal{X}$ minus parameter $\gamma>0$ times the variance of the embedded perturbation of the ground-truth backdoor patterns in the poisoned training subset, $(x+\delta_{{\rm B},x},t)\in\mathcal{X}_{\rm B}\subset\mathcal{X}$:
$-\sum_{(x,y)\in\mathcal{X}}\log\hat{p}(y|x)$
$-~{}\frac{\gamma}{|\mathcal{X}_{\rm B}|}\sum_{(x+\delta_{{\rm B},x},t)\in\mathcal{X}_{\rm B}}\|f(x+\delta_{\rm B,x})-f(x)-\mu_{\rm B}\|^{2},$
where
$\mu_{\rm B}=\frac{1}{|\mathcal{X}_{\rm B}|}\sum_{(x+\delta_{{\rm B},x},t)\in\mathcal{X}_{\rm B}}(f(x+\delta_{\rm B,x})-f(x)).$
Based on the previous equation, we launched an adaptive attack on a Resnet-18 model trained on the CIFAR-10 dataset. The backdoor trigger was jointly optimized with the Resnet-18 model over 100 epochs, using a learning rate of 0.001. In one experiment, the adaptive attack parameter $\gamma$ was set to 1 and the poisoning rate was 0.5 (50%). Consensus maximization for the adaptive attack was conducted in the penultimate layer of the Resnet-18 model. Under the adaptive attack, a high poisoning rate was needed to achieve a reasonable attack success rate (ASR), which in this case was 0.75. The clean-data accuracy (ACC) of the poisoned model was 0.83.
Subsequently, we implemented CEPA using the $\|\mu\|$ criterion in the embedded feature space of the $5^{\rm th}$ layer of the poisoned Resnet-18 model. The backdoor was successfully detected with a MAD anomaly score of 46.36 for the true target class.